DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, January 3 Padelford C-36 3:45 PM

Bi-Lipschitz equivalence and Jacobians

Dmitri Burago (Penn State)

We construct a separated net in the plane that is not bi-Lipschitz equivalent to a periodic lattice. The key ingredient is a construction of a continuous function that is not a Jacobian of a smooth (or even Lipschitz) homeomorphism - answering a problem by Moser. There are some further developments and intriguing open questions.

This is a joint work with B. Kleiner.